# QSO MUSEUM II: Search for extended Ly emission around 8 quasar pairs

### v1

@@
QSO MUSEUM II: Search for extended Ly emission around 8 quasar pairs
Extended Ly emission is routinely found around single quasars across cosmic time. However, few studies have investigated how such emission changes in fields with physically associated quasar pairs, which should reside in dense environments and are predicted to be linked through intergalactic filaments. We present VLT/MUSE snapshot observations (45 minutes/source) to unveil extended Ly emission on scales of the circumgalactic medium (CGM) around the largest sample of physically associated quasar pairs to date, encompassing 8 pairs (14 observed quasars) at with -band magnitude between 18 and 22.75, corresponding to absolute magnitudes between -29.6 and -24.9. The pairs are either at close (50-100 kpc, 5 pairs) or wide (450-500 kpc, 3 pairs) angular separation, and have velocity differences of km s-1. We detect extended emission around 12 of the 14 targeted quasars and investigate the luminosity, size, kinematics and morphology of these Ly nebulae. On average, they span about 90 kpc and are 2.8 erg s-1 bright. Irrespective of the quasars’ projected distance, the nebulae often (45 %) extend toward the other quasar in the pair, leading to asymmetric emission whose flux-weighted centroid is at an offset position from any quasar location. We show that large nebulae are preferentially aligned with the large-scale structure as traced by the direction between the two quasars, and conclude that the cool gas (104 K) in the CGM traces well the direction of cosmic web filaments. Additionally, the radial profile of the Ly surface brightness around quasar pairs can be described by a power law with a shallower slope () with respect to single quasars (), indicative of increased CGM densities out to large radii and/or enhanced contribution from the intergalactic medium (IGM) due to the dense environments expected around quasar pairs. The sample presented in this study contains excellent targets for ultra-deep observations to directly study filamentary IGM structures in emission. This work demonstrates that a large snapshot survey of quasar pairs will pave the way to the statistical direct study of the IGM.
@@
In the current paradigm of galaxy evolution, the circumgalactic medium (CGM) plays a major role in regulating star formation and active galactic nuclei (AGN) activity. Loosely defined to be the gas virially bound to a galaxy, but beyond the visible stellar disk (Tumlinson et al., 2017), the CGM is believed to be heavily influenced by galactic outflows, heating and chemically enriching the gas reservoir, as well as inflows of cool gas (104 K) from the intergalactic medium (IGM). While this cold accretion from the IGM filamentary gas structures, also known as the cosmic web, onto the CGM, is ubiquitously seen in state-of-the-art cosmological simulations (e.g., Dubois et al., 2014; Schaye et al., 2015; Springel et al., 2018), it is difficult to directly observe due to the expected low densities.
@@
Constraints on CGM properties, for instance gas density, temperature, metallicity and kinematics, have been achieved through line-of-sight absorption studies (e.g., Hennawi et al., 2006; Prochaska et al., 2013; Rudie et al., 2019; Péroux & Howk, 2020; Donahue & Voit, 2022; Faucher-Giguère & Oh, 2023) or by directly observing CGM emission with narrowband filters or long-slit spectra (e.g., Hu & Cowie 1987; Heckman et al. 1991; Weidinger et al. 2005; Hennawi & Prochaska 2013). However, the new generation of sensitive integral field unit (IFU) spectrographs like the Multi Unit Spectroscopic Explorer (MUSE; Bacon et al. 2010) on the Very Large Telescope (VLT) and the Keck Cosmic Web Imager (KCWI; Morrissey et al. 2012) have made it possible to routinely observe the emission of the cool CGM gas around individual galaxies by utilizing, for example, bright nebular lines like Ly. Most commonly, high- quasars () are targeted for these studies (e.g., Borisova et al., 2016; Arrigoni Battaia et al., 2019a; Farina et al., 2019; Cai et al., 2019; O’Sullivan et al., 2020; Mackenzie et al., 2021; Fossati et al., 2021), often unveiling large portions of the host galaxies’ CGM with the bulk of the Ly emission extending for kpc. Indeed quasars are estimated to reside in relatively massive halos across cosmic times ( M⊙; Shen et al. 2007; White et al. 2012; Eftekharzadeh et al. 2015; Fossati et al. 2021; Pizzati et al. 2024; Costa 2024), resulting in a virial radius of kpc at , close to their cosmic noon.
@@
While these AGNs produce copious amounts of ionizing photons, greatly outshining their host galaxies, the exact balance between powering mechanisms of their Ly nebulae remains disputed. The suggested possibilities include the combination of recombination radiation after the ionization of the material by the quasar ionizing photons (Cantalupo et al., 2005; Kollmeier et al., 2010; Costa et al., 2022), resonant scattering of Ly photons from the quasar (Cantalupo et al., 2014; Costa et al., 2022), shocks (Taniguchi & Shioya, 2000), gravitational cooling radiation (Haiman et al., 2000; Dijkstra et al., 2006). Additionally, in-situ scattering of Ly photons is likely important in shaping the nebulae morphology (Costa et al., 2022; Arrigoni Battaia et al., 2023b). Notwithstanding these uncertainties, both current observations and simulations emphasize the main role of the quasar radiation in powering the extended emission (Costa et al., 2022; González Lobos et al., 2023; Langen et al., 2023; Obreja et al., 2024).
@@
The most extreme Ly nebulae known, extending up to 500 projected kpc and therefore exceeding the predicted virial radius for quasar halos, have been detected around systems hosting multiple quasars: the Slug nebula, situated between a radio-loud and a radio-quiet quasar at redshift , extends for 460 kpc (Cantalupo et al., 2014); Hennawi et al. (2015) reported the detection of four quasars at embedded in a 310 kpc long nebula; Arrigoni Battaia et al. (2018) presented a Ly nebula extending over 297 kpc and associated with one bright and two faint quasars at ; and Cai et al. (2018) detected a 232 kpc wide nebula surrounding a quasar pair. This phenomenon might be explained by the overdensity of UV luminous galaxies able to power nebular emission and the surplus of gas fed into the system through tidal interactions during close encounters, which, in turn, could also fuel the AGN activity.
@@
Subsequently, targeted quasar pair fields often show remarkable extended emission. Arrigoni Battaia et al. (2019b) discovered a Ly nebula stretching between two quasars with a projected separation of kpc. Li et al. (2023) targeted four pairs consisting of one bright ( mag) and one faint quasar at with the Palomar Cosmic Web Imager (PCWI; Matuszewski et al. 2010) and found a maximum projected extent upward of 200 kpc in three of them, indicative of cool gas outside of the CGM being illuminated. A quasar pair at with a projected separation of 500 kpc observed for 40 hours as part of the MUSE Ultra Deep Field (MUDF) program hosts multiple Ly nebulae extending up to 100 kpc and preferentially along the connecting line between the quasars (Lusso et al., 2019). This is interpreted as possible evidence of a large scale gaseous filament connecting the two galaxies.
@@
Given the aforementioned expected quasar halo mass of 1012.5 M⊙, quasar pairs could indeed be suitable tracers of filament orientations. In the MilleniumTNG simulation (Pakmor et al. 2023), the mean halo mass at the nodes of the cosmic web is 1012.47 M⊙ at redshift 3, with almost all halos above 1013 M⊙ being identified as a node (Galárraga-Espinosa et al., 2024). Per definition, nodes indicate the endpoint of cosmic web filaments and therefore two neighboring nodes are expected to be connected to each other by a gaseous filament.
@@
Here, we extend these previous works by searching for extended nebular emission around a sample of 8 quasar pairs observed with VLT/MUSE. Our work has been conducted in the framework of the survey Quasar Snapshot Observations with MUse: Search for Extended Ultraviolet eMission (QSO MUSEUM; Arrigoni Battaia et al. 2019a), using a very similar strategy. The paper is structured as follows. Section 2 describes the observations and the data reduction. In Sect. 3 we explain how the data are analyzed to unveil the extended Ly emission. Sections 4 and 5 present the results of this work and discuss them in comparison to previous studies. Finally, Sect. 6 summarizes our findings. Throughout this paper, we assume a flat CDM cosmology with km s-1 Mpc-1, and . At the mean redshift of our sample, , one arcsecond corresponds to roughly 7.7 kpc. Reported quasar magnitudes are in the AB system and absolute magnitudes are K-corrected relative to redshift 2 (Ross et al., 2013).
[…]
| 1.2 | 00:01:40.5984 | 07:09:47.8149 | 22.75 | -24.93 | 3.238 | ||||
| 2.2 | 00:18:08.0959 | 16:12:50.8420 | 22.13 | -25.42 | 3.126 | ||||
| 3.2 | 02:40:05.7477 | -00:39:14.0616 | 22.39 | -25.16 | 3.120 | ||||
[…]
| 6.2 | 10:21:16.9849 | 11:12:27.5716 | 20.28 | -28.01 | 3.815 | ||||
| 7.2 | 10:40:49.1050 | 01:00:33.0900 | 20.85 | -26.59 | 3.026 | ||||
| 8.2 | 11:35:02.5085 | -02:21:20.1432 | 22.06 | -25.36 | 3.009 |
@@ 2 Observations and data reduction
The sample was selected from the SDSS DR12 quasar catalog (Pâris et al., 2017) complemented by AGN searches at high- (e.g., Bielby et al. 2013) as described in Arrigoni Battaia et al. (2019b). In brief, to ensure physical association of the quasar pairs, targets are selected to be within of each other (or about km s-1) based on their systemic redshift provided by the aforementioned works, and have an angular separation of 1 arcmin ( kpc). Moreover, the redshift of the quasar pairs was restricted to the range . The lower limit ensures that the Ly line is within the wavelength range of MUSE, while the upper limit prevents it from being heavily contaminated by sky emission lines. Although the sample selected in this way comprised 17 pairs covering a wide range of projected distances between the two quasars, only close ( 50-100 kpc) and wide ( 450 kpc) pairs have been observed due to weather conditions. In the two widest pairs, only one quasar is targeted in each system due to the large angular separation. One quasar pair from the selected and observed sample (ID 8.1 and 8.2) has already been published in Arrigoni Battaia et al. (2019b). We included it here to increase the sample size, and re-analyzed it with the same procedures for comparison. The full spectra of the quasars are shown in Fig. 1. If the quasar is not in the FoV of the MUSE observations, its SDSS spectra is instead shown. In the VLA FIRST survey (Becker et al., 1994), a radio source is detected 0.8 away from quasar 5.1 with . No other quasar from this sample is detected in VLA FIRST. Figure 2 displays the redshift and magnitude of quasars in our sample in comparison to previous surveys of both single quasar nebulae and quasar pairs.
@@
The sample presented in this paper has been observed from November 2017 to March 2018 as part of the program 0100.A-0045(A) (PI: F. Arrigoni Battaia) in service mode with the MUSE instrument in Wide Field Mode on the VLT telescope YEPUN. Quasar properties and information about the observations are summarized in Table 1. The observations have been conducted at an average seeing of 1.16” with clear weather conditions or thin cloud coverage. For each target, three exposures with 880 seconds of on source time each were taken, rotated by 90 degrees with respect to each other and a few arcseconds dithering pattern. The only exception is quasar pair 7, for which only 2 exposures are available.
@@
Each cube covers a wavelength range of 4750 - 9350 with a spectral resolution at the mean wavelength of Ly of and at a spectral sampling of 1.25 and a field of view of approximately 1x1 with a spatial sampling of 0.2. The data has been reduced using the MUSE pipeline v2.8.7 (Weilbacher et al., 2014, 2016, 2020) as described in Farina et al. (2019) and González Lobos et al. (2023), applying bias-subtraction, flat-fields, twilight and illumination correction, sky-subtraction, wavelength and flux calibration. After running the pipeline, there are still residual sky emission lines affecting the data. Therefore, we apply an additional skyline subtraction using the Zurich Atmospheric Purge tool (ZAP, Soto et al. 2016), minimizing both the residuals of sky emission lines and the influence of this procedure on the flux of emission at the expected wavelength and position of the Ly glow. Subsequently, artifacts due to the edges of the MUSE integral field units are masked via visual inspection and the offset between individual exposures is determined by centroid fitting of a bright point source in the FoV. Then, the science exposures and variance cubes for each object are median combined. To determine a more accurate estimate of the noise by taking into account the correlated noise introduced by the pipeline interpolations, the associated variance cubes are rescaled layer by layer to the RMS spectrum of a sigma-clipped background region of the science cubes as usually done in the literature (e.g., Borisova et al. 2016; Bacon et al. 2017). This final variance cube is subsequently used to calculate the surface brightness limit (SBL) and errors. The average SBL in a 30 pseudo narrow-band (PNB) image centered on Ly and after masking continuum sources via sigma-clipping is in a 1 arcsec2 aperture, while the average SBL in the same aperture and in the respective central slice is . This is comparable, but lower than the layerwise SBL reported in Arrigoni Battaia et al. (2019a) () and Borisova et al. (2016) (). Both surveys have comparable observational setups, but include more targets at , where sky emission lines can lower the sensitivity at the Ly line.
@@ 3 Data analysis
To reveal the extended emission, the quasar point spread function (PSF) has to be subtracted from the science data cube as it outshines the extended emission (Møller et al., 2000; Christensen et al., 2006; Husemann et al., 2014). Here we use the same empirical method described extensively in González Lobos et al. (2023). Specifically, the size of the PSF and therefore the required subtraction window are dependent on the seeing, which is estimated from the quasar continuum emission. Narrowbands of 25 width are created and a 2D Moffat function is fitted onto the quasar profile while conservatively excluding spectral regions where extended emission might occur, that is, slices around the quasar emission lines typically spanning around 500 Å. Extrapolated to the Ly wavelength at the systemic redshift of the respective quasar, the Moffat FWHM provides a good estimate of the seeing of the observations. In order to obtain more robust results, we perform this estimation on the brighter quasar in the field, although the results are consistent for all pairs if performed individually. Additionally, we determine the seeing at the wavelength of Ly using a star in the FoV, if available. Especially for faint quasars, the host galaxy contribution can be high enough that the quasar no longer appears as a point source, leading to an overestimation of the seeing. However, the seeing determined from a star is on average only 0.3 % smaller than the one determined from the respective quasar, with the biggest deviation being 6 %.
@@
In previous works the PSF subtraction has been performed in circular windows with a diameter ranging from 6 to 10 times the seeing depending on the luminosity of the quasar (Borisova et al. 2016; Mackenzie et al. 2021), with brighter quasars requiring a larger subtraction region. Given the diversity of luminosities in our sample, we accordingly vary the window size between objects in a similar range as previous studies (circular aperture with diameter 4 to 9 times the seeing), aiming to have virtually no contribution from the quasar PSF to the flux value at the edge of the circular region. If the chosen window size is too big, oversubtraction can occur, and sometimes lead to spurious emission detection in the PSF and the continuum subtracted data cube if there are absorption features in the Ly line of the quasar.
@@
An empirical PSF is constructed layerwise by creating narrowbands in the subtraction window and scaling it to the flux value in the central 1x1. In this normalization area, residual values are very error-prone and the pixels are therefore masked for subsequent analysis (as usually done in similar analyses, e.g., Borisova et al. 2016). To avoid including possible extended nebular emission in the PSF construction, wavelength slices containing quasar emission lines are skipped when constructing narrowbands and instead, the next narrowband redward of the line is used to create the PSF. This is especially important for the Ly line as the narrowband at shorter wavelengths than the line can be heavily affected by absorption. The width of narrowbands has to be varied depending on the magnitude of the quasar in order to achieve sufficient signal-to-noise ratios to constrain the PSF shape, with a typical width of 350 channels or 437.5 for this sample. The empirical PSF constructed in this way is subtracted from the respective quasar in the data cube to reveal the large scale emission around it. Subtraction of two PSFs is not necessary in every field: In the quasar pairs 4 and 5, only the brighter quasar is within the MUSE FoV due to the large angular separation, and quasar 1.2 is very dim with an -band magnitude of 24.6 at the time of discovery with VLT/VIMOS (Bielby et al., 2013) and is detected as a faint continuum source in the MUSE observation (-band 22.75). Therefore, only one PSF was subtracted in these fields.
@@
Also, two systems (quasar 5.1 and 3.2) have “non-standard” PSF subtractions. In quasar 5.1, continuum sources very close to the PSF disturb the subtraction and have therefore been masked before constructing the empirical PSF. Examining the nature of these sources will be the focus of a companion paper (Herwig et al. in prep). The PSF of quasar 3.2 has a very asymmetric spectral shape, leading to severe oversubtraction in one half of the subtraction window when using our standard approach, with a mean value within a 30 narrowband around the Ly peak equivalent to -2.3, and undersubtraction in the other half. This is likely due to a very bright narrow nebular line compared to the fainter and broader Ly from the quasar. Therefore, we mask the quasar spectrum in the slices where nebular emission is suspected and replace the values with an interpolation. During the construction, the PSF is then scaled by the masked spectrum. The mask encompasses 12 slices, or 15 , and is chosen to be the narrowest mask capable of removing the oversubtraction.
@@
Lastly, we utilize the ZAP package to estimate the continuum using the function contsubfits which applies a median filter. The obtained continuum cube is then subtracted from the PSF subtracted datacube obtaining a datacube which should only contain nebular emission lines and noise. In the subsequent analysis, continuum sources, excluding the quasars, are masked.

### v2

# QSO MUSEUM. II. Search for extended Ly emission around eight quasar pairs

@@
QSO MUSEUM. II. Search for extended Ly emission around eight quasar pairs
Extended Ly emission is routinely found around single quasars across cosmic time. However, few studies have investigated how such emission changes in fields with physically associated quasar pairs, which should reside in dense environments and are predicted to be linked through intergalactic filaments. We present VLT/MUSE snapshot observations (45 minutes/source) to unveil extended Ly emission on scales of the circumgalactic medium (CGM) around the largest sample of physically associated quasar pairs to date, encompassing eight pairs (14 observed quasars) at with an -band magnitude between 18 and 22.75, corresponding to absolute magnitudes between -29.6 and -24.9. The pairs are either at close (50-100 kpc, five pairs) or wide (450-500 kpc, three pairs) angular separation and have velocity differences of km s-1. We detected extended emission around 12 of the 14 targeted quasars and investigated the luminosity, size, kinematics, and morphology of these Ly nebulae. On average, they span about 90 kpc and are 2.8 erg s-1 bright. Irrespective of the quasars’ projected distance, the nebulae often (45 %) extend toward the other quasar in the pair, leading to asymmetric emission whose flux-weighted centroid is at an offset position from any quasar location. We show that large nebulae are preferentially aligned with the large-scale structure, as traced by the direction between the two quasars, and conclude that the cool gas (104 K) in the CGM traces well the direction of cosmic web filaments. Additionally, the radial profile of the Ly surface brightness around quasar pairs can be described by a power law with a shallower slope () with respect to single quasars (), indicative of increased CGM densities out to large radii and/or an enhanced contribution from the intergalactic medium (IGM) due to the dense environments expected around quasar pairs. The sample presented in this study contains excellent targets for ultra-deep observations to directly study filamentary IGM structures in emission. This work demonstrates that a large snapshot survey of quasar pairs will pave the way to direct statistical study of the IGM.
@@
In the current paradigm of galaxy evolution, the circumgalactic medium (CGM) plays a major role in regulating star formation and the activity of active galactic nuclei (AGNs). Loosely defined to be the gas virially bound to a galaxy but beyond the visible stellar disk (Tumlinson et al., 2017), the CGM is believed to be heavily influenced by galactic outflows, heating and chemically enriching the gas reservoir, as well as inflows of cool gas (104 K) from the intergalactic medium (IGM). While this cold accretion from the IGM filamentary gas structures, also known as the cosmic web, onto the CGM is ubiquitously seen in state-of-the-art cosmological simulations (e.g., Dubois et al., 2014; Schaye et al., 2015; Springel et al., 2018), it is difficult to directly observe due to the expected low densities.
@@
Constraints on CGM properties (e.g., gas density, temperature, metallicity, and kinematics) have been achieved through line-of-sight absorption studies (e.g., Hennawi et al., 2006; Prochaska et al., 2013; Rudie et al., 2019; Péroux & Howk, 2020; Donahue & Voit, 2022; Faucher-Giguère & Oh, 2023) or by directly observing CGM emission with narrowband filters or long-slit spectra (e.g., Hu & Cowie 1987; Heckman et al. 1991; Weidinger et al. 2005; Hennawi & Prochaska 2013). However, the new generation of sensitive integral field unit (IFU) spectrographs such as the Multi Unit Spectroscopic Explorer (MUSE; Bacon et al. 2010) on the Very Large Telescope (VLT) and the Keck Cosmic Web Imager (KCWI; Morrissey et al. 2012) have made it possible to routinely observe the emission of the cool CGM gas around individual galaxies by utilizing, for example, bright nebular lines such as Ly. Most commonly, high- quasars () are targeted for such studies (e.g., Borisova et al., 2016; Arrigoni Battaia et al., 2019a; Farina et al., 2019; Cai et al., 2019; O’Sullivan et al., 2020; Mackenzie et al., 2021; Fossati et al., 2021), often unveiling large portions of the host galaxies’ CGM and with the bulk of the Ly emission extending for kpc. Indeed, quasars are estimated to reside in relatively massive halos across cosmic times ( M⊙; Shen et al. 2007; White et al. 2012; Eftekharzadeh et al. 2015; Fossati et al. 2021; Pizzati et al. 2024; Costa 2024), resulting in a virial radius of kpc at , close to their cosmic noon.
@@
While these AGNs produce copious amounts of ionizing photons, greatly outshining their host galaxies, the exact balance between the powering mechanisms of their Ly nebulae remains disputed. The suggested possibilities include the combination of recombination radiation after the ionization of the material by the quasar ionizing photons (Cantalupo et al., 2005; Kollmeier et al., 2010; Costa et al., 2022), resonant scattering of Ly photons from the quasar (Cantalupo et al., 2014; Costa et al., 2022), shocks (Taniguchi & Shioya, 2000), gravitational cooling radiation (Haiman et al., 2000; Dijkstra et al., 2006). Additionally, in situ scattering of Ly photons is likely important in shaping the nebulae morphology (Costa et al., 2022; Arrigoni Battaia et al., 2023b). Notwithstanding these uncertainties, both current observations and simulations emphasize the main role of the quasar radiation in powering the extended emission (Costa et al., 2022; González Lobos et al., 2023; Langen et al., 2023; Obreja et al., 2024).
@@
The most extreme Ly nebulae known, commonly referred to Enormous Ly nebulae (ELANe), extend up to 500 projected kiloparsec and therefore exceed the predicted virial radius for quasar halos. They have been detected around systems hosting multiple quasars: the Slug nebula, situated between a radio-loud and a radio-quiet quasar at redshift , which extends for 460 kpc (Cantalupo et al., 2014); Hennawi et al. (2015) reported the detection of four quasars at embedded in a 310 kpc long nebula; Arrigoni Battaia et al. (2018) presented a Ly nebula extending over 297 kpc and associated with one bright and two faint quasars at ; and Cai et al. (2018) detected a 232 kpc wide nebula surrounding a quasar pair. This phenomenon might be explained by the overdensity of UV luminous galaxies able to power nebular emission and the surplus of gas fed into the system through tidal interactions during close encounters, which in turn could also fuel AGN activity.
@@
Subsequently, targeted quasar pair fields often show remarkable extended emission. Arrigoni Battaia et al. (2019b) discovered a Ly nebula stretching between two quasars with a projected separation of kpc. Li et al. (2023) targeted four pairs consisting of one bright ( mag) and one faint quasar at with the Palomar Cosmic Web Imager (PCWI; Matuszewski et al. 2010) and found a maximum projected extent upward of 200 kpc in three of them, indicative of cool gas outside the CGM being illuminated. A quasar pair at with a projected separation of 500 kpc observed for 40 hours as part of the MUSE Ultra Deep Field (MUDF) program hosts multiple Ly nebulae extending up to 100 kpc and preferentially along the connecting line between the quasars (Lusso et al., 2019). This is interpreted as possible evidence of a large-scale gaseous filament connecting the two galaxies.
@@
Given the aforementioned expected quasar halo mass of 1012.5 M⊙, quasar pairs could indeed be suitable tracers of filament orientations. In the MilleniumTNG simulation (Pakmor et al. 2023), the mean halo mass at the nodes of the cosmic web is 1012.47 M⊙ at redshift three, with almost all halos above 1013 M⊙ being identified as a node (Galárraga-Espinosa et al., 2024). Per definition, nodes indicate the endpoint of cosmic web filaments, and therefore, two neighboring nodes are expected to be connected to each other by a gaseous filament.
@@
In this paper, we extend the research of previous works by searching for extended nebular emission around a sample of eight quasar pairs observed with VLT/MUSE. Our work has been conducted in the framework of the survey Quasar Snapshot Observations with MUse: Search for Extended Ultraviolet eMission (QSO MUSEUM; Arrigoni Battaia et al. 2019a), using a very similar strategy. The paper is structured as follows. Section 2 describes the observations and the data reduction. In Sect. 3, we explain how the data are analyzed to unveil the extended Ly emission. Sections 4 and 5 present the results of this work and discuss them in comparison to previous studies. Finally, Sect. 6 summarizes our findings. Throughout this paper, we assume a flat CDM cosmology with km s-1 Mpc-1, , and . At the mean redshift of our sample, , one arcsecond corresponds to roughly 7.7 kpc. Reported quasar magnitudes are in the AB system, and absolute magnitudes are K-corrected relative to redshift two (Ross et al., 2013).
@@ 2 Observations and data reduction
The sample was selected from the SDSS DR12 quasar catalog (Pâris et al., 2017) complemented by AGN searches at high- (e.g., Bielby et al. 2013) as described in Arrigoni Battaia et al. (2019b). In brief, to ensure the physical association of the quasar pairs, targets were selected to be within of each other (or about km s-1) based on their systemic redshift provided by the aforementioned works and to have an angular separation of 1 arcmin ( kpc). Moreover, the redshift of the quasar pairs was restricted to the range . The lower limit ensures that the Ly line is within the wavelength range of MUSE, while the upper limit prevents it from being heavily contaminated by sky emission lines. Although the sample selected in this way comprised 17 pairs covering a wide range of projected distances between the two quasars, only close ( 50-100 kpc) and wide ( 450 kpc) pairs were observed due to weather conditions. In the two widest pairs, only one quasar was targeted in each system due to the large angular separation. One quasar pair from the selected and observed sample (ID 8.1 and 8.2) has already been published in Arrigoni Battaia et al. (2019b). We included it here to increase the sample size, and re-analyzed it with the same procedures for comparison. The full spectra of the quasars are shown in Fig. 1. If the quasar is not in the field of view (FoV) of the MUSE observations, its SDSS spectra is instead shown. In the VLA FIRST survey (Becker et al., 1994), a radio source is detected 0.8 away from quasar 5.1 with . No other quasar from this sample is detected in VLA FIRST. Figure 2 displays the redshift and magnitude of quasars in our sample in comparison to previous surveys of both single quasar nebulae and quasar pairs.
@@
The sample presented in this paper was observed from November 2017 to March 2018 as part of the program 0100.A-0045(A) (PI: F. Arrigoni Battaia) in service mode with the MUSE instrument in Wide Field Mode on the VLT telescope YEPUN. Quasar properties and information about the observations are summarized in Table 4. The observations were conducted at an average seeing of 1.16” (or kpc at the mean redshift of the sample) with clear weather conditions or thin cloud coverage. For each target, three exposures with 880 seconds of on source time each were taken, rotated by 90 degrees with respect to each other and a few arcseconds dithering pattern. The only exception is quasar pair 7, for which only two exposures are available.
@@
Each cube covers a wavelength range of 4750 - 9350 with a spectral resolution at the mean wavelength of Ly of (corresponding to or 165 km s-1) and at a spectral sampling of 1.25 and a field of view of approximately 1x1 with a spatial sampling of 0.2. The data was reduced using the MUSE pipeline v2.8.7 (Weilbacher et al., 2014, 2016, 2020) as described in Farina et al. (2019) and González Lobos et al. (2023), applying bias-subtraction, flat-fields, twilight and illumination correction, sky-subtraction, wavelength and flux calibration. After running the pipeline, there are still residual sky emission lines affecting the data. Therefore, we applied an additional skyline subtraction using the Zurich Atmospheric Purge tool (ZAP, Soto et al. 2016), minimizing both the residuals of sky emission lines and the influence of this procedure on the flux of emission at the expected wavelength and position of the Ly glow. Subsequently, we masked artifacts due to the edges of the MUSE integral field units via visual inspection and determined the offset between individual exposures by centroid fitting of a bright point source in the FoV. Then, we median combined the science exposures and variance cubes for each object. To determine a more accurate estimate of the noise by taking into account the correlated noise introduced by the pipeline interpolations, we rescaled the associated variance cubes layer by layer to the RMS spectrum of a sigma-clipped background region of the science cubes as usually done in the literature (e.g., Borisova et al. 2016; Bacon et al. 2017). This final variance cube was subsequently used to calculate the surface brightness limit (SBL) and errors. The average SBL in a 30 pseudo narrowband (PNB) image centered on Ly and after masking continuum sources via sigma-clipping is in a 1 arcsec2 aperture, while the average SBL in the same aperture and in the respective central slice is . This is comparable, but lower than the layerwise SBL reported in Arrigoni Battaia et al. (2019a) () and Borisova et al. (2016) (). Both surveys have comparable observational setups, but include more targets at , where sky emission lines can lower the sensitivity at the Ly line.
@@ 3 Data analysis
To reveal the extended emission, the quasar point spread function (PSF) has to be subtracted from the science data cube as it outshines the extended emission (Møller et al., 2000; Christensen et al., 2006; Husemann et al., 2014). Here we used the same empirical method described extensively in González Lobos et al. (2023). Specifically, the size of the PSF and therefore the required subtraction window are dependent on the seeing, which we estimated from the quasar continuum emission. We created narrowbands of 25 width and fitted a 2D Moffat function onto the quasar profile while conservatively excluding spectral regions where extended emission might occur, that is, slices around the quasar emission lines typically spanning around 500 Å. Extrapolated to the Ly wavelength at the systemic redshift of the respective quasar, the Moffat FWHM provides a good estimate of the seeing of the observations. In order to obtain more robust results, we performed this estimation on the brighter quasar in the field, although the results are consistent for all pairs if performed individually. Additionally, we determined the seeing at the wavelength of Ly using a star in the FoV, if available. Especially for faint quasars, the host galaxy contribution can be high enough that the quasar no longer appears as a point source, leading to an overestimation of the seeing. However, the seeing determined from a star is on average only 0.3 % smaller than the one determined from the respective quasar, with the biggest deviation being 6%.
@@
In previous works the PSF subtraction was performed in circular windows with a diameter ranging from 6 to 10 times the seeing depending on the luminosity of the quasar (Borisova et al. 2016; Mackenzie et al. 2021), with brighter quasars requiring a larger subtraction region. Given the diversity of luminosities in our sample, we accordingly varied the window size between objects in a similar range as previous studies (circular aperture with diameter 4 to 9 times the seeing), aiming to have virtually no contribution from the quasar PSF to the flux value at the edge of the circular region. If the chosen window size is too big, oversubtraction can occur, and sometimes lead to spurious emission detection in the PSF and the continuum subtracted data cube if there are absorption features in the Ly line of the quasar.
@@
We constructed an empirical PSF layerwise by creating narrowbands in the subtraction window and scaling it to the flux value in the central 1x1. In this normalization area, residual values are very error-prone and we therefore masked the pixels for subsequent analysis (as usually done in similar analyses, e.g., Borisova et al. 2016). To avoid including possible extended nebular emission in the PSF construction, wavelength slices containing quasar emission lines were skipped when constructing narrowbands and instead, the next narrowband redward of the line was used to create the PSF. This is especially important for the Ly line as the narrowband at shorter wavelengths than the line can be heavily affected by absorption. The width of narrowbands has to be varied depending on the magnitude of the quasar in order to achieve sufficient signal-to-noise ratios to constrain the PSF shape, with a typical width of 350 channels or 437.5 for this sample. We subtracted the empirical PSF constructed in this way from the respective quasar in the data cube to reveal the large scale emission around it. Subtraction of two PSFs was not necessary in every field: In the quasar pairs 4 and 5, only the brighter quasar is within the MUSE FoV due to the large angular separation, and quasar 1.2 is very dim with an -band magnitude of 24.6 at the time of discovery with VLT/VIMOS (Bielby et al., 2013) and is detected as a faint continuum source in the MUSE observation (-band 22.75). Therefore, we subtracted only one PSF in these fields.
@@
Also, two systems (quasar 5.1 and 3.2) required “non-standard” PSF subtractions. In quasar 5.1, continuum sources very close to the PSF disturb the subtraction and we therefore masked them before constructing the empirical PSF. Examining the nature of these sources will be the focus of a companion paper Herwig et al. 2024.The PSF of quasar 3.2 has a very asymmetric spectral shape, leading to severe oversubtraction in one half of the subtraction window when using our standard approach, with a mean value within a 30 narrowband around the Ly peak equivalent to -2.3, and undersubtraction in the other half. This is likely due to a very bright narrow nebular line compared to the fainter and broader Ly from the quasar. Therefore, we masked the quasar spectrum in the slices where nebular emission is suspected and replaced the values with an interpolation. During the construction, we then scaled the PSF by the masked spectrum. The mask encompasses 12 slices, or 15 , and was chosen to be the narrowest mask capable of removing the oversubtraction.
@@
Lastly, we utilized the ZAP package to estimate the continuum using the function contsubfits, which applies a median filter. The obtained continuum cube was then subtracted from the PSF-subtracted data cube, yielding a data cube that should only contain nebular emission lines and noise. In the subsequent analysis, continuum sources, excluding the quasars, were masked.